OpenStudy (anonymous):

can u explain how to workout the answer to this question 1 A laser emits light of wavelength 600nm. What is the distance, expressed as a number of wavelengths, travelled by the light in one second? A 5 × 108 B 5 × 1011 C 5 × 1014 D 5 × 1017

OpenStudy (anonymous):

d = distance D = 600nm c = 3x10^8 m/s t = 1 s v=d/t d=v*t d=(3*10^8 m/s)*(1 s) d=3*10^8 m n: How many D's "fit" inside d (d expressed as n*D). d=n*D n=(3*10^8 m)/(600*10^-9 m) n= 0,005*10^17 n= 5*10^14 C) 5x10^14